If a function is defined as f(x,n) mod n. Determine the

i.     Domain of f

ii.   Range of f

iii.   G(g(g(g(7)))) if g (n) = f(209, n).


1
Expert's answer
2021-09-24T06:41:02-0400

Part i

The mod function we can only apply on integers

So, Domain "\{f(x,n)\}= Set \space of \space integers = \{...-2,-1,0,1,2...\}"


Part ii

Range of f(x,n)= {0,1,2,...n-1}

Because x and n = remainder when xn divided by nx


Part iii

"g(7)=f(209,7)= 209 mod 7= 6\\
g(g(7))=g(6)=f(209,6)= 5\\
g(g(g(7)))=f(209,5)= 4\\
g(g(g(g(7))))=f(209,4)= 5\\"